# Deza Cayley graphs from difference sets

### v1

@@ Abstract.
In this note, we provide several constructions of Deza Cayley graphs over groups having a generalized dihedral subgroup. These constructions are based on a usage of (relative) difference sets.
@@
The author was supported by the grant of The Natural Science Foundation of Hebei Province (project No. A2023205045)
@@ 1. Introduction
A regular graph [^1] Throughout the paper, by a graph we mean a simple undirected graph without loops. $\Gamma$ is said to be strongly regular if there exist nonnegative integers $\lambda$ and $\mu$ such that every two adjacent vertices of $\Gamma$ have exactly $\lambda$ common neighbors, whereas every two distinct nonadjacent vertices of $\Gamma$ have exactly $\mu$ common neighbors. Nowadays, strongly regular graphs are realized as one of the crucial objects in algebraic combinatorics. For a background of strongly regular graphs, we refer the readers to the monograph [4].
@@
The numbers $(v,k,b,a)$, where $v$ is the number of vertices of $\Gamma$ and $k$ is the degree of each vertex, are called the parameters of $\Gamma$. We may always assume that $a\leq b$. Clearly, if $a>0$ and $b>0$, then $\Gamma$ has diameter $2$. A Deza graph is called a strictly Deza graph if it is not strongly regular and has diameter $2$. Deza graphs have been actively studied during the last years. For more details on Deza graphs and a recent progress in their studying, we refer the readers to the survey paper [12]. Computational results on Deza graphs can be found in [10].
@@
One of the main problems concerned with Deza graphs is the problem of constructing new families of them. In the present paper, we are interested in this problem for Deza Cayley graphs. By a Cayley graph $\Gamma=\operatorname{Cay}(G,S)$ over a finite group $G$ with a non-empty inverse-closed identity-free connection set $S\subseteq G$, we mean the graph with vertex set $G$ and edge set $E=\{(g,xg):~{}x\in X,~{}g\in G\}$. All Deza Cayley graphs with at most $60$ vertices were enumerated in [11]. Several results on Deza Cayley graphs over cyclic groups are given in [3] and [12, Section 2], whereas several results on Deza Cayley graphs over dihedral groups are given in [18]. A construction of a Deza Cayley graph with the largest possible association scheme and the smallest possible automorphism group can be found in [5].
The main goal of this paper is to provide new constructions of Deza Cayley graphs. The key ingredient for this is a usage of (relative) difference sets.
@@
A subset $R$ of $G$ is called a relative difference set (RDS for short) in $G$ if there exist a normal subgroup $N$ of $G$ and a nonnegative integer $\lambda$ such that every element of $G\setminus N$ has exactly $\lambda$ representations in the form $g_{1}g_{2}^{-1}$, where $g_{1},g_{2}\in R$, and no non-identity element of $N$ has such a representation.
The subgroup $N$ and the numbers $(m,n,k,\lambda)$, where $m=|G:N|$, $n=|N|$, and $k=|R|$, are called the forbidden subgroup and parameters of $R$, respectively. The RDS $R$ is said to be nontrivial if $22\leq|R|\leq|G|-$ 2. The latter implies that $\lambda>0$ and $m>1$. It is easy to check that $R$ contains at most one element from each $N$-coset. If $R$ contains exactly one element from each $N$-coset, i.e. $R$ is a transversal for $N$ in $G$, then $R$ is said to be semiregular. In this case, $m=k=\lambda n$. If $N$ is trivial, then $R$ is called a difference set (DS for short). If $R$ is a DS, then the parameters of $R$ are written as $(v,k,\lambda)$, where $v=|G|$.
An RDS is a special case of a divisible difference set which can be considered as a Cayley object, i.e. a combinatorial object having a regular subgroup in its automorphism group, in the class of divisible designs. For a background of DSs and RDSs, we refer the readers to [2, Section VI] and [16], respectively.
Let us finish the introduction with a brief outline of the paper. We construct a subset $S$ of a generalized dihedral group $G$ from an (R)DS $R$ in the abelian subgroup of $G$ of index $2$ and establish necessary and sufficient conditions for $S$ to be a connection set of a Deza Cayley graph in terms of the parameters of $R$ (Propositions 3.1 and 3.5). Further, we show that there exist (R)DSs satisfying the above conditions (Lemmas 3.2 and 3.6) and construct Deza Cayley graphs (Corollaries 3.3 and 3.7). Applying several operations to the latter graphs, we obtain more Deza Cayley graphs (Corollaries 3.9 and 3.10). The main tool used in the proofs of the above results is computations in the integer group ring of $G$. A necessary background of group rings is given in Section $2$.
2. Group ring
Let $G$ be a finite group and $\mathbb{Z}G$ the integer group ring. The identity element and the set of all nonidentity elements of $G$ are denoted by $e$ and $G^{\#}$, respectively. A product of two elements $x=\sum_{g\in G}x_{g}g,y=\sum_{g\in G}y_{g}g\in\mathbb{Z}G$ will be written as $x\cdot y$. If $X\subseteq G$, then the set $\{x^{-1}:x\in X\}$ and the element $\sum\limits_{x\in X}{x}$ of the group ring $\mathbb{Z}G$ are denoted by $X^{-1}$ and $\underline{X}$, respectively.
It is easy to verify that
Let $N$ be a normal subgroup of $G$, $m=|G:N|$, and $n=|N|$. It follows easily from the definition of RDS that a subset $R$ of $G$ is an RDS with forbidden subgroup $N$ and parameters $(m,n,k,\lambda)$, where $k=|R|$ and $\lambda\geq 0$, if and only if
The following lemma provides an easy criterion for a Cayley graph to be a Deza graph in terms of the integer group ring. It can be found, e.g, in [3, Lemma $5.2$].
Lemma 2.1.
$\underline{S}^{2}=ke+a\underline{S_{a}}+b\underline{S_{b}},$
3. Constructions
In this section, we provide the constructions of Deza Cayley graphs based on a usage of DSs and RDSs and two operations which allow to construct new Deza Cayley graphs from the above ones. Throughout the section, $A$ is an abelian group and $G=A\rtimes\langle b\rangle$, where $|b|=2$ and $a^{b}=a^{-1}$ for every $a\in A$, i.e. $G$ is a generalized dihedral group associated with $A$.
3.1. Construction from DSs
The following statement in case of a dihedral group was proved in [18, Lemma $3.1$]. However, the proof in case of a generalized dihedral group is exactly the same.
Proposition 3.1.
$k=\frac{2v-1-\sqrt{8v-7}}{2}.$
In this case, $\Gamma(R)$ is a strictly Deza graph and has parameters $(2v,v-1+k,2k,2(k-1))$.
It is well-known that if $R$ is a DS in $A$ with parameters $(v,k,\lambda)$, then $(A\setminus R)$ is a DS in $A$ with parameters $(v,v-k,v-2k+\lambda)$.
@@ 3.1. Construction from DSs
According to [2, Table A.3.1], there exist DSs with parameters $(7,3,1)$, $(11,5,2)$, $(16,6,2)$, and $(37,9,2)$ in abelian groups. The first of them and the complements to the other ones are the required in the lemma DSs. ∎
The DS with parameters $(7,3,1)$ is of Singer type, whereas the DSs with parameters $(11,6,3)$, $(16,10,6)$, and $(37,28,21)$ are the complements to the DSs of Paley, Hadamard, and biquadric residue types, respectively (see [2, Appendix, Section 3]). Clearly, a group of order $7$, $11$, or $37$ is necessary cyclic. Due to [2, Table A.3.1], a DS with parameters $(16,10,6)$ exists in each group of order $16$ except for the cyclic one.
It is easy to verify that each DS $R$ from Lemma 3.2 satisfies the condition $k=\frac{2v-1-\sqrt{8v-7}}{2}$ from Proposition 3.1 and hence $\Gamma(R)$ is a strictly Deza Cayley graph. So we obtain the next corollary.
Corollary 3.3.
There exist strictly Deza Cayley graphs with parameters
$(14,9,6,4),~{}(22,16,12,10),~{}\text{and}~{}(74,64,56,54)$
over dihedral groups and a strictly Deza Cayley graph with parameters
$(32,25,20,18)$
over a generalized dihedral group associated with any noncyclic group of order $16$.
All the graphs from Corollary 3.3 with at most $60$ vertices appear in the list from [11]. We do not know whether there exist other DSs satisfying the condition $k=\frac{2v-1-\sqrt{8v-7}}{2}$. Checking using the database of DSs [9] implies that there is no other such a DS with $v\leq 10000$.
3.2. Constructions from RDSs
We start with an easy lemma on RDSs.
Lemma 3.4.
Let $R$ be an RDS in $A$ with a forbidden subgroup $A_{0}$. Then $A_{0}R=A$ if and only if $R$ is semiregular.
Proof.
Recall that $R$ contains at most one element from each $A_{0}$-coset. So the equality $A_{0}R=A$ is equivalent to the fact that $R$ is a transversal for $A_{0}$ in $A$ which means that $R$ is semiregular. ∎
Proposition 3.5.
  1. (1)
    The graph $\Gamma_{1}(R)=\operatorname{Cay}(G,S_{1})$, where $S_{1}=Rb\cup A_{0}^{\#}$, is a Deza graph if and only if
    $(n,\lambda)=(2,2)~{}\text{or}~{}(n,\lambda)=(4,2).$
    In this case, $\Gamma_{1}(R)$ has parameters
    $(2mn,k+n-1,2,0)$
    and it is of diameter $2$ if and only if $R$ is semiregular. If the latter is the case, then $\Gamma_{1}(R)$ is strongly regular if and only if $(n,\lambda)=(2,2)$.
  2. (2)
    The graph $\Gamma_{2}(R)=\operatorname{Cay}(G,S_{2})$, where $S_{2}=G^{\#}\setminus Rb$, is a Deza graph if and only if
    $\lambda=2.$
    In this case, $\Gamma_{2}(R)$ is a strictly Deza graph and has parameters
    $(2mn,2mn-k-1,2mn-2k,2mn-2k-2).$
Proof.
Observe that $\lambda>0$ because $R$ is nontrivial. Let us prove Statement $(1)$. Clearly, $|G|=2|A|=2mn$ and $|Rb\cup A_{0}^{\#}|=|Rb|+|A_{0}|-1=k+n-1$ and hence $\Gamma_{1}(R)$ is a $(k+n-1)$-regular graph on $2mn$ vertices. A straightforward computation in the group ring of $G$ using the equality $a^{b}=a^{-1}$ for every $a\in A$ and Eqs. (2) and (3) implies that
Since $R$ contains at most one element from each $A_{0}$-coset, every element of $A$ enters the element $\underline{A_{0}}^{\#}\cdot\underline{R}$ with coefficient 00 or $1$. Moreover, every element of $R$ enters $\underline{A_{0}}^{\#}\cdot\underline{R}$ with coefficient 00 because $e\notin A_{0}^{\#}$. From Lemma 2.1 and Eq. (4) it follows that $\Gamma_{1}(R)$ is a Deza graph if and only if
|$|\{n-2,\lambda,0,2\}|\leq 2.$
Since $\lambda>0$, the latter is equivalent to $(n,\lambda)\in\{(2,2),(4,2)\}$. If the latter holds, then $\Gamma_{1}(R)$ has the required parameters by Lemma 2.1 and Eq. (4). Observe that $\Gamma_{1}(R)$ has diameter $2$ if and only if every element of
$G^{\#}\setminus S_{1}=(A\setminus A_{0})\cup(A\setminus R)b$
$A\setminus R\subseteq A_{0}^{\#}R.$
Now let us prove Statement $(2)$. In this case, $|G^{\#}\setminus Rb|=|G|-1-|Rb|=2mn-k-1$. So $\Gamma_{2}(R)$ is a $(2mn-k-1)$-regular graph on $2mn$ vertices. One can compute using the equality $a^{b}=a^{-1}$ for every $a\in A$ and Eqs. (1), (2), and (3) that
|$|\{2(mn-k-1),2(mn-k),2(mn-k-1)+\lambda\}|\leq 2.$
Since $\lambda>0$, the latter is equivalent to $\lambda=2$. The remaining part of Statement $(2)$ is a consequence of Lemma 2.1 and Eq. (5). ∎
The lemma below provides families of RDSs which can be used for the constructions from Proposition 3.5.
@@ 3.2. Constructions from RDSs
Statement $(1)$ follows from [6, Theorem $1$] applied to $a=k$ and $b=k-1$. Statements $(2)$ and $(3)$ are special cases of [1, Theorem $1.2$] when $d=2$ and $n=\frac{q-1}{2}$ and [15, Theorem 2.1] when $p=2$ and $c=1$, respectively. Finally, Statement $(4)$ is taken from [14] (see also [1, Result $3$]). ∎
Corollary 3.7.
The following statements hold.
@@
    There exist a strictly Deza Cayley graph with parameters
    $(64,11,2,0)$
    over the generalized dihedral group associated with the group $\mathbb{Z}_{4}^{2}\times\mathbb{Z}_{2}$ and non-strictly Deza Cayley graphs with parameters
    $(24,6,2,0)~{}\text{and}~{}(80,12,2,0)$
    over the dihedral groups of orders $24$ and $80$, respectively.
@@
    There exist strictly Deza Cayley graphs with parameters
    $(2^{2k},2^{2k}-2^{k}-1,2^{2k}-2^{k+1},2^{2k}-2^{k+1}-2),$
    $(q^{2}-1,q^{2}-q-2,q^{2}-2q-1,q^{2}-2q-3),$

### v2

@@ Abstract.
In this note, we provide several constructions of Deza Cayley graphs over groups having a generalized dihedral subgroup. These constructions are based on usage of (relative) difference sets.
@@
The author was supported by the state contract of the Sobolev Institute of Mathematics (project number FWNF-2022-0017)
@@ 1. Introduction
A regular graph [^1] Throughout the paper, by a graph we mean a simple undirected graph without loops. $\Gamma$ is said to be strongly regular if there exist nonnegative integers $\lambda$ and $\mu$ such that every two adjacent vertices of $\Gamma$ have exactly $\lambda$ common neighbors, whereas every two distinct nonadjacent vertices of $\Gamma$ have exactly $\mu$ common neighbors. Nowadays, strongly regular graphs are realized as one of the keynote objects in algebraic combinatorics. For a background of strongly regular graphs, we refer the readers to the monograph [6].
@@
The numbers $(v,k,b,a)$, where $v$ is the number of vertices of $\Gamma$ and $k$ is the degree of each vertex, are called the parameters of $\Gamma$. We may always assume that $a\leq b$. Clearly, if $a>0$ and $b>0$, then $\Gamma$ has diameter $2$. A Deza graph is called a strictly Deza graph if it is not strongly regular and has diameter $2$.
Deza graphs attract attention of several mathematicians. This class of graphs has been actively studied during the last years. Deza graphs with special parameters were investigated in [13, 20, 21]. Some results on spectra of Deza graphs were obtained in [1, 2]. A vertex connectivity of Deza graphs was studied in [11, 14]. The class of Deza graphs with strongly regular children was a research object in [19]. For more details on Deza graphs and a recent progress in their studying, we refer the readers to the survey paper [17]. Computational results on Deza graphs can be found in [15].
@@
In the present paper, we are interested in Deza Cayley graphs. By a Cayley graph $\Gamma=\operatorname{Cay}(G,S)$ over a finite group $G$ with a non-empty inverse-closed identity-free connection set $S\subseteq G$, we mean the graph with vertex set $G$ and edge set $E=\{(g,xg):~x\in X,~g\in G\}$. All Deza Cayley graphs with at most $60$ vertices were enumerated in [16]. Several results on Deza Cayley graphs over cyclic groups are given in [5] and [17, Section 2], whereas several results on Deza Cayley graphs over dihedral groups are given in [27]. A construction of a Deza Cayley graph with the largest possible association scheme and the smallest possible automorphism group can be found in [7].
One of the main problems concerned with Deza graphs is the problem of constructing new families of them. In this paper, we provide several constructions of Deza Cayley graphs. Recall that a generalized dihedral group associated with an abelian group $A$ is defined to be a semidirect product of $A$ and a cyclic group of order $2$ whose nontrivial element inverses every element of $A$. The main result of the paper is the theorem below stating an existence of two new infinite families of strictly Deza Cayley graphs over generalized dihedral groups.
Theorem 1.1.
There exist strictly Deza Cayley graphs with parameters
$(2^{2k},2^{2k}-2^{k}-1,2^{2k}-2^{k+1},2^{2k}-2^{k+1}-2)$
and
$(q^{2}-1,q^{2}-q-2,q^{2}-2q-1,q^{2}-2q-3),$
over the generalized dihedral groups associated with the groups $\mathbb{Z}_{4}^{k-1}\times\mathbb{Z}_{2}$ and $\mathbb{Z}_{\frac{q^{2}-1}{2}}$, respectively, where $k\geq 3$ and $q\geq 5$ is an odd prime power.
As we can check using the database of results on Deza graphs [24], the graphs from Theorem 1.1 are previously unknown except for the graphs with parameters $(24,18,14,12)$ and $(48,40,34,32)$ from the second family for $q=5$ and $q=7$ which appear among small Deza Cayley graphs from [16]. The constructions of Deza Cayley graphs from Theorem 1.1 are based on usage of relative difference sets. Our technique enables us to find some more sporadic Deza Cayley graphs.
Theorem 1.2.
There exists a Deza Cayley graph with parameters $(v,k,b,a)$ from the first column of Table $1$ over a generalized dihedral group associated with an abelian group $A$ from the second column of Table $1$.
$(v,k,b,a)$ AA strictly
$(14,9,6,4)$ $\mathbb{Z}_{7}$ yes
$(16,5,2,0)$ $\mathbb{Z}_{4}\times\mathbb{Z}_{2}$, $\mathbb{Z}_{2}^{3}$ no
$(16,11,8,6)$ $\mathbb{Z}_{4}\times\mathbb{Z}_{2}$, $\mathbb{Z}_{2}^{3}$ yes
$(22,16,12,10)$ $\mathbb{Z}_{11}$ yes
$(24,6,2,0)$ $\mathbb{Z}_{12}$ no
$(32,25,20,18)$ $\mathbb{Z}_{8}\times\mathbb{Z}_{2}$, $\mathbb{Z}_{4}^{2}$, $\mathbb{Z}_{4}\times\mathbb{Z}_{2}^{2}$, $\mathbb{Z}_{2}^{4}$ yes
$(64,11,2,0)$ $\mathbb{Z}_{4}^{2}\times\mathbb{Z}_{2}$ yes
$(74,64,56,54)$ $\mathbb{Z}_{37}$ yes
$(80,12,2,0)$ $\mathbb{Z}_{40}$ no
$(252,235,220,218)$ $\mathbb{Z}_{126}$ yes
Table 1. Sporadic Deza Cayley graphs
As we are able to verify, the last four graphs from Table $1$ are previously unknown, whereas the other ones can be found in the list of small Deza Cayley graphs from [16]. The graph with parameters $(16,5,2,0)$ is isomorphic to the strongly regular van Lint-Schrijver graph (see [6, p. 371]).
Let us finish the introduction with a brief outline of the paper. Deza graphs from Theorems 1.1 and 1.2 are constructed from difference sets and relative difference sets (see Propositions 4.1 and 4.2). A necessary information and families of (relative) difference sets used for constructing Deza Cayley graphs are given in Section $2$. One of the main tools used in the proof of Proposition 4.2 is computations in the integer group ring of $G$. A short background of group rings is given in Section $3$. In Section $4$, we provide constructions of Deza Cayley graphs and prove Theorems 1.1 and 1.2. We observe in Section $5$ that parameters of the graphs from Theorems 1.1 and 1.2 satisfy some arithmetic conditions which allows to construct from these graphs more Deza graphs using some operations on graphs. Several concluding remarks are given in Section $6$.
Acknowledgement: The author would like to thank the anonymous referee for comments which help to improve the text substantially.
2. Difference sets
In this section, we give necessary information on difference sets. At first, let us recall the definition of a relative difference set.
@@ 2. Difference sets
A subset $R$ of $G$ is called a relative difference set (RDS for short) in $G$ if there exist a subgroup $N$ of $G$ and a nonnegative integer $\lambda$ such that every element of $G\setminus N$ has exactly $\lambda$ representations in the form $g_{1}g_{2}^{-1}$, where $g_{1},g_{2}\in R$, and no non-identity element of $N$ has such a representation.
The subgroup $N$ and the numbers $(m,n,k,\lambda)$, where $m=|G:N|$, $n=|N|$, and $k=|R|$, are called the forbidden subgroup and parameters of $R$, respectively. The RDS $R$ is said to be nontrivial if $2\leq|R|\leq|G|-2$. The latter implies that $\lambda>0$ and $m>1$. It is easy to check that $R$ contains at most one element from each right $N$-coset. If $R$ contains exactly one element from each right $N$-coset, i.e. $R$ is a right transversal for $N$ in $G$, then $R$ is said to be semiregular. In this case, $m=k=\lambda n$.
If $N$ is trivial, then $R$ is called a difference set (DS for short). If $R$ is a DS, then the parameters of $R$ are written as $(v,k,\lambda)$, where $v=|G|$. It is well-known that if $R$ is a DS in $A$ with parameters $(v,k,\lambda)$, then $(A\setminus R)$ is a DS in $A$ with parameters
$(v,v-k,v-2k+\lambda).$
An RDS is a special case of a divisible difference set which can be considered as a Cayley object, i.e. a combinatorial object having a regular subgroup in its automorphism group, in the class of divisible designs. For a background of DSs and RDSs, we refer the readers to [4, Section VI] and [25], respectively.
Below, we provide families of DSs and RDSs which will be used for constructing Deza Cayley graphs.
@@
According to [4, Table A.3.1], there exist DSs with parameters $(7,3,1)$, $(11,5,2)$, $(16,6,2)$, and $(37,9,2)$ in abelian groups. The first of them and the complements to the other ones are the required DSs in the lemma. ∎
@@
Statement $(1)$ follows from [8, Theorem 11] applied to $a=k$ and b=k1b=k-1. Statements $(2)$ and $(3)$ are special cases of [3, Theorem $1.2$] when $d=2$ and $n=\frac{q-1}{2}$ and [23, Theorem 2.1] when $p=2$ and $c=1$, respectively. Finally, Statement $(4)$ is taken from [22] (see also [3, Result 33]). ∎
3. Group ring
Let $G$ be a finite group and $\mathbb{Z}G$ the integer group ring. The identity element and the set of all nonidentity elements of $G$ are denoted by $e$ and $G^{\#}$, respectively. A product in the group ring of two elements $x=\sum_{g\in G}x_{g}g,y=\sum_{g\in G}y_{g}g\in\mathbb{Z}G$ will be written as $x\cdot y$. If $X\subseteq G$, then the set $\{x^{-1}:x\in X\}$ and the element $\sum\limits_{x\in X}{x}$ of the group ring $\mathbb{Z}G$ are denoted by $X^{-1}$ and $\underline{X}$, respectively.
It is easy to verify that
Let $N$ be a subgroup of $G$, $m=|G:N|$, and $n=|N|$. It follows easily from the definition of RDS that a subset $R$ of $G$ is an RDS with forbidden subgroup $N$ and parameters $(m,n,k,\lambda)$, where $k=|R|$ and $\lambda\geq 0$, if and only if
Lemma 3.1.
Let $R$ be an RDS in a group $G$ with a forbidden subgroup $N$. Then $\underline{N}\cdot\underline{R}=G$ if and only if $R$ is semiregular.
Proof.
Recall that $R$ contains at most one element from each right $N$-coset. So the equality $\underline{N}\cdot\underline{R}=G$ is equivalent to the fact that $R$ is a transversal for $N$ in $G$ which means that $R$ is semiregular. ∎
The following lemma provides an easy criterion for a Cayley graph to be a Deza graph in terms of the integer group ring. It can be found, e.g, in [5, Lemma $5.2$].
Lemma 3.2.
$\underline{S}^{2}=ke+a\underline{S_{a}}+b\underline{S_{b}},$
4. Constructions
In this section, we provide the constructions of Deza Cayley graphs based on usage of DSs and RDSs. Throughout the section, $A$ is an abelian group and $G=A\rtimes\langle b\rangle$, where $|b|=2$ and $a^{b}=a^{-1}$ for every $a\in A$, i.e. $G$ is a generalized dihedral group associated with $A$.
The following statement in case of a dihedral group was proved in [27, Lemma $3.1$]. However, the proof in case of a generalized dihedral group is exactly the same.
Proposition 4.1.
$k=\frac{2v-1-\sqrt{8v-7}}{2}.$
In this case, $\Gamma(R)$ is a strictly Deza graph and has parameters (2v,v1+k,2k,2(k1))(2v,v-1+k,2k,2(k-1)).
Proposition 4.2.
@@ 4. Constructions
    The graph $\Gamma_{1}(R)=\operatorname{Cay}(G,S_{1})$, where $S_{1}=Rb\cup A_{0}^{\#}$, is a Deza graph if and only if (n,λ)=(2,2) or (n,λ)=(4,2). In this case, $\Gamma_{1}(R)$ has parameters (2mn,k+n-1,2,0) and it is of diameter $2$ if and only if $R$ is semiregular. If the latter is the case, then $\Gamma_{1}(R)$ is strongly regular if and only if $(n,\lambda)=(2,2)$.
@@
    The graph $\Gamma_{2}(R)=\operatorname{Cay}(G,S_{2})$, where $S_{2}=G^{\#}\setminus Rb$, is a Deza graph if and only if λ=2. In this case, $\Gamma_{2}(R)$ is a strictly Deza graph and has parameters (2mn,2mn-k-1,2mn-2k,2mn-2k-2).
Proof.
Observe that $\lambda>0$ because $R$ is nontrivial. Let us prove Statement $(1)$. Clearly, $|G|=2|A|=2mn$ and $|Rb\cup A_{0}^{\#}|=|Rb|+|A_{0}|-1=k+n-1$ and hence $\Gamma_{1}(R)$ is a (k+n1)(k+n-1)-regular graph on $2mn$ vertices. A straightforward computation in the group ring of $G$ using the equality $a^{b}=a^{-1}$ for every $a\in A$ and Eqs. (2) and (3) implies that
Since $R$ contains at most one element from each $A_{0}$-coset, every element of $A$ enters the element $\underline{A_{0}}^{\#}\cdot\underline{R}$ with coefficient 0 or $1$. Moreover, every element of $R$ enters $\underline{A_{0}}^{\#}\cdot\underline{R}$ with coefficient 0 because $e\notin A_{0}^{\#}$. From Lemma 3.2 and Eq. (4) it follows that $\Gamma_{1}(R)$ is a Deza graph if and only if
$|\{n-2,\lambda,0,2\}|\leq 2$
or, equivalently,
$\{n-2,\lambda\}\subseteq\{0,2\}.$
Since $\lambda>0$, we conclude that $\lambda=2$ and hence $(n,\lambda)\in\{(2,2),(4,2)\}$. If the latter holds, then $\Gamma_{1}(R)$ has the required parameters by Lemma 3.2 and Eq. (4). Observe that $\Gamma_{1}(R)$ has diameter $2$ if and only if every element of
$G^{\#}\setminus S_{1}=(A\setminus A_{0})\cup(A\setminus R)b$
$A\setminus R\subseteq A_{0}^{\#}R.$
Now let us prove Statement $(2)$. In this case, $|G^{\#}\setminus Rb|=|G|-1-|Rb|=2mn-k-1$. So $\Gamma_{2}(R)$ is a (2mnk1)(2mn-k-1)-regular graph on $2mn$ vertices. One can compute using the equality $a^{b}=a^{-1}$ for every $a\in A$ and Eqs. (1), (2), and (3) that
$|\{2(mn-k-1),2(mn-k),2(mn-k-1)+\lambda\}|\leq 2,$
or, equivalently,
$2(mn-k-1)+\lambda\in\{2(mn-k-1),2(mn-k)\}.$
Since $\lambda>0$, we obtain $\lambda=2$. The remaining part of Statement $(2)$ is a consequence of Lemma 3.2 and Eq. (5). ∎
Proof of Theorems 1.1 and 1.2.
Theorem 1.1 is a consequence of Statement $(2)$ of Proposition 4.2 applied to the RDSs from Statements $(1)$ and $(2)$ of Lemma 2.2.
It is easy to verify that the parameters of each DS from Lemma 2.1 satisfy the condition $k=\frac{2v-1-\sqrt{8v-7}}{2}$ and hence applying Proposition 4.1 to these DSs, we obtain the strictly Deza Cayley graphs with parameters $(14,9,6,4)$, $(22,16,12,10)$, $(32,25,20,18)$, and $(74,64,56,54)$ from Theorem 1.2. Other graphs from Theorem 1.2 can be obtained by applying Statement $(1)$ of Proposition 4.2 to the RDSs from Statement $(1)$ with $k=3$, Statement $(2)$ with $q=5$ and $q=9$, and Statement $(3)$ of Lemma 2.2 and applying of Statement $(2)$ of Proposition 4.2 to the RDSs from Statements $(3)$ and $(4)$ of Lemma 2.2. ∎
5. More Deza graphs
In this section, we construct more Deza Cayley graphs using some operations. The first of the operations is the lexicographic product of graphs (see, e.g., [10, p. 398]). Recall that the lexicographic product $\Gamma_{1}[\Gamma_{2}]$ of graphs $\Gamma_{1}=(V_{1},E_{1})$ and $\Gamma_{2}=(V_{2},E_{2})$ is defined to be the graph with vertex set $V=V_{1}\times V_{2}$ and edge set $E$ defined as follows:
$((v_{1},v_{2}),(u_{1},u_{2}))\in E~\text{if and only if}~(v_{1},u_{1})\in E_{1}~\text{or}~v_{1}=u_{1}~\text{and}~(v_{2},u_{2})\in E_{2}.$
Recall that the Cartesian product $\Gamma_{1}\times\Gamma_{2}$ of $\Gamma_{1}$ and $\Gamma_{2}$ is defined to be the graph with vertex set $V=V_{1}\times V_{2}$ and edge set $E$ defined as follows:
$((v_{1},v_{2}),(u_{1},u_{2}))\in E~\text{if and only if}~v_{1}=u_{1}~\text{and}~(v_{2},u_{2})\in E_{2}~\text{or}~v_{2}=u_{2}~\text{and}~(v_{1},u_{1})\in E_{1}.$
The clique on $m$ vertices is denoted by $K_{m}$ and the complement to the graph $\Gamma$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$. The second operation is defined as follows:
$\Gamma\mapsto\mkern 1.5mu\overline{\mkern-1.5mu\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu\times K_{2}\mkern-1.5mu}\mkern 1.5mu.$
Using the definitions, it is easy to verify that $\mkern 1.5mu\overline{\mkern-1.5mu\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu\times K_{2}\mkern-1.5mu}\mkern 1.5mu$ is isomorphic to a disjoint union of two copies of $\Gamma$ in which each vertex from the one copy is adjacent to each vertex from the other one except its own copy.
Lemma 5.1.
Let $\Gamma$ be a strictly Deza graph with parameters (v,k,2kv+2,2kv)(v,k,2k-v+2,2k-v). Then $K_{m}[\Gamma]$, $m\geq 1$, and $\mkern 1.5mu\overline{\mkern-1.5mu\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu\times K_{2}\mkern-1.5mu}\mkern 1.5mu$ are strictly Deza graphs with parameters
(mv,k+(m1)v,(m2)v+2k+2,(m2)v+2k)(mv,k+(m-1)v,(m-2)v+2k+2,(m-2)v+2k)